Help please, don't have time. due 11:59pm 02/29. Please answer all. thanks!4. Compute the equivalent units of production for materials. Equivalent units of production for materialsRequired information The Foundational 15 (LO4-1, LO4-2, LO4-3, LO4-4, LO4-5) [The following information applies to the questiRequired: 1. Prepare the journal entries to record the raw materials used in production and the direct labor cost incurred. (2. Prepare the journal entry to record the overhead cost applied to production. (If no entry is required for a transaction/ev3. How many units were completed and transferred to finished goods during the period? Completed and transferred out

4. Compute the equivalent units of production for materials. Equivalent units of production for materials5. Compute the equivalent units of production for conversion. Equivalent units of production6. What is the cost of beginning work in process inventory plus the cost added during the period for materials? Total cost of7. What is the cost of beginning work in process inventory plus the cost added during the period for conversion? Total cost o8. What is the cost per equivalent unit for materials? (Round your answer to 2 decimal places.) Cost per equivalent unit for9. What is the cost per equivalent unit for conversion? (Round your answer to 2 decimal places.) Cost per equivalent unit10. What is the cost of ending work in process inventory for materials? (Round your intermediate calculations to 2 places.) C
Answer #1
Journal entries
TR General Journal Debit Credit
1) Work in process-Mixing 120,205
Raw materials inventory 120,205
2) Work in process-Mixing 98,000
Manufacturing overhead 98,000
3) Completed & transferred out 34,500
4) Equivalent units of production for materials 42,700
5) Equivalent units of production 38,600
6) Total cost of materials 138,775
7) Total cost of conversion 194,930
8) Cost per equivalent units pf materials 3.25
9) Cost per equivalent units 5.05
10) Cost of ending wip inventory for materials 26650

Working

Beginning work in process inventory 5,100 Completed & transferred 34,500
started during the year 37,600 Ending work in process 8,200
units to be accounted 42,700 units accounted for 42,700
Equivalent units
% mat Materials % CC Conversion
Completed & transferred 34,500 100% 34,500 100% 34,500
ending work in process 8,200 100% 8,200 50% 4100
42,700 42,700 38,600
cost per unit
Materials CC Total
Beginning work in process 18,570 16,430 35,000
cost added during the year 120,205 178500 298,705
total cost 138,775 194,930 333,705
equivalent units 42,700 38,600
cost per unti 3.25 5.05
cost of ending work in process inventory.
materials 8,200 * 3.25 = 26650
conversion 4,100 * 5.05 = 20705
cost of ending work in process inventory. 47355
cost of goods transferred to finished goods
materials 34,500 * 3.25 = 112125
conversion 34,500 * 5.05 = 174225
cost of ending work in process inventory. 286350